### v2

@@ 2.3 Jet injection
where $t_{s}$ represents the typical sound crossing time scale in the ambient gas, which is calculated within a characteristic feedback radius $R_{\text{fb}}$. $R_{\text{fb}}$ is the radius within which the initial thermal energy of the ambient gas is equal to the jet outburst energy (Duan & Guo 2020). As $R_{\text{fb}}$  is determined by the initial gas distribution and jet injection energy, both of which are fixed in our simulations, it remains the same for all our runs. In the context of this work, we have $t_{s}\approx 50\text{ Myr}$ and $P_{\text{fb}}\approx 1.3\times 10^{44}\text{ erg s}^{-1}$. The physical meaning of $P_{\text{fb}}$ lies in the fact that if the jet power, $P_{\text{jet}}$, significantly exceeds $P_{\text{fb}}$, the jet will undergo noticeable shock dissipation. Furthermore, taking into account the Eddington luminosity of the black hole $L_{\text{Edd}}\approx 6.71\times 10^{46}\text{ erg s}^{-1}$, the ratio of jet power to Eddington luminosity is $\lambda_{\text{jet}}=P_{\text{jet}}/L_{\text{Edd}}$. For the case where $P_{\text{jet}}=P_{\text{fb}}$, this ratio yields $\lambda_{\text{jet}}\approx 2\times 10^{-3}$.
@@
Figure 2: Scales of the radio sources, $d_{25}$ at the radio power $10^{25}\rm{WHz^{-1}}$ (left) and the corresponding times (right), varying with the jet powers, $P_{\rm jet}$, in simulations with diverse jet parameters, as shown in Table 1. The dots in the right panel represent the time $t_{25}$ when the radio power of the source declines to $10^{25}\rm{WHz^{-1}}$. The horizontal lines in the right panel denote the jet duration $t_{\rm jet}$ of the corresponding powers. The triangles denote $t_{\rm jet}+t_{\rm syn}(t_{\rm jet})$ (upward) and $t_{\rm jet}+t_{\rm syn}(t_{\rm jet})$ (downward), where $t_{\rm syn}(t)$ is the typical synchrotron cooling time at 144(1+z)MHz for the volume averaged magnetic field strength corresponding to the time $t$, calculated using Eq. 39. The jet powers are normalized by the typical feedback power (bottom) and Eddington luminosity (top), as discussed in Sect. 2.3. The stars are color-coded to represent distinct fractions of energy components and the magnetic structure parameter $\alpha_{p}$, as indicated in the plot.
@@ 3 Results
It is also interesting to see whether the ”lower power-larger bubble” effect exists when considering the same radio power for different sources. In the left panel of Fig. 2, we present the scales of the radio sources varying with jet power, specifically at the instant when the sources have declined to radio power (luminosity) of $10^{25}\rm{WHz^{-1}}$ at 144 (1+z) $\rm{MHz}$ for redshift z=0.3. The methodology for calculating radio emission is detailed in Appendix C. Notably, we can see a trend that the lower power jets result in larger radio sources. The corresponding time $t_{25}$ for the radio power of $10^{25}\rm{WHz^{-1}}$ in different runs are shown in the right panel of Fig. 2 (dots), and we can see they all occur after the jets are switched off ($t_{25}>t_{\rm jet}$ ). We also calculate the typical synchrotron cooling times for the toroidal magnetic field dominant runs at times $t=t_{\rm jet}$) and $t=t_{25}$. We can observe that, if the electrons cannot be adequately accelerated after the jets are switched off, they will quickly cool down, particularly for the high-power jets (upward triangles in the right panel of Fig. 2). This means that, when considering synchrotron cooling, the high-power jets will actually result in further smaller radio sources for the same radio power.
@@ 3.3 Radio power of the simulated sources
On the other hand, although jets with significantly lower power can generate much larger radio sources, they may be too faint to observe. As illustrated in Fig. 3, the evolution of the lowest-power jet (run t1200fm8, represented by the black solid line) barely reaches the lower boundary of the observational region. This result implies that the jet powers of these radio sources lie approximately within the range of $10^{-4}-10^{-3}L_{\rm Edd}$ (the corresponding ratios of the jet powers to the Eddington luminosity are shown in Fig. 2). Furthermore, the jet active time of run t1200fm8 is approximately the typical synchrotron cooling time (Eq. 39) of the electrons in the radio lobes, thus the outer part of the lobes might be radio dark if new acceleration of electrons is not adequate. Considering the very low density and pressure of the extreme low-power jets, the mixing between ejecta and surrounding medium might even let them fail to form low density bubbles or observable radio lobes.
@@ 4.1 Energy transport processes
It should be significant to see the energy transport processes in different runs before we discuss in detail the production of large radio sources. We now focus on two sets of runs: the runs with the same energy fractions ($f_{\rm m}=0.8$, $f_{\rm th}=0.05$, $\alpha_{\rm p}=0.3$) but different powers and the runs with the same power ($t_{\rm jet}=50\ \rm Myr$) but different energy fractions or magnetic field structure. The upper panels in Fig. 5 show the evolution of the distance that the jetted bubble traveled (half of the dimension of radio sources) in these two sets of runs. The evolution of the total energy residing in jetted bubbles is also shown in the lower panels of 5. The dotted vertical lines show when the same total energy is injected in different runs. We can see the scales of the jetted bubbles generally shows positive correlation with the total energy left in the bubbles particularly for the jets with the same energy compositions but different active times (powers). Notably, the high-power jet (run t5fm8) expels energy more rapidly upon jet cessation, leaving less energy within the bubble, which impedes the growth of the radio sources. Actually the most high-power jets can transmit energy outward through more potent shocks compared to low-power jets, which rely on less potent processes such as sound waves and slow expansions (Tang & Churazov 2017; Bambic & Reynolds 2019; Duan & Guo 2020).
To show more details of energy transport, we calculate the power of the energy outflows at the surfaces of jetted bubbles in different runs and the main results are shown in Figs. 6 and 11. As shown in the upper left panel of Fig. 6, we can see that the higher power jets can produce much stronger energy transport in the center of the cluster halo, while the lower power jets undergo more energy outflow flux at the outer region. On the other hand, to check how the energy flows out, we decompose the energy outflow into three forms: (1) the forward pushing energy outflow for the energy flowing out along the $z+$ direction; (2) the transverse expanding energy outflow for the energy flowing out along the $r+$ direction; and (3) the back flow disturbing energy outflow, which is actually negligible in our simulations. The fractions of the power of the forward pushing energy outflow and the transverse expanding energy outflow are shown in the lower panels in the Fig. 6. We can see that the energy is mainly transported out forward in the inner region and transversely in the outer region for almost all the magnetic dominant jets; kinetic and thermal jets, however, mainly transport energy out transversely in the inner region. We also show the 2D plots of the evolution of total energy outflow flux for run t200fm8 in Fig. 10 and the power of different kind of energy outflow in Fig. 11, which may be of interest to some readers.
@@ 4.2 Production of large radio sources
To gain some details about the production of large radio sources, we then examine the evolution of the widths of the jetted bubbles and show the results in the upper panels of Fig. 7. We can see that, within the central region of the halos (for example, r¡100kpc), the widths of the bubbles generally exhibit an anti-correlation with the scale of the radio sources ultimately displayed in the simulations (see Figs. 1 and 5). Furthermore, we check the volume-averaged internal pressure (magnetic pressure plus thermal pressure) within the bubbles and present the results in the lower panels of Fig. 7. We can see that internal pressures of bubbles show a positive correlation with the widths of bubbles in the innermost region of the halos. The dotted vertical lines show where the correlations have just begun to develop. This phenomenon should be due to the fact that higher internal pressures within bubbles will trigger more prominent transverse expansion, resulting in larger cross sections. The larger cross sections will make the jets much less penetrating and easier to be decelerated by the dense gas in the center of the halos. In the upper left panel of Fig. 7, we can observe that higher power jets generally produce wider bubbles at the center of the halos. Subsequently, there will be stronger interactions between the higher power jets and the surrounding gas, causing them to transfer more energy than lower power jets in the center of the halos, as discussed above.
The differences between the jets in run t50fm8 and run t50fm8bp3 are very subtle, as they are injected with the same energy fractions and power but have different magnetic structures. As the emerging jet is far from equilibrium with the surrounding gas, it will undergo rapid changes in its properties. Before the identified low-density bubbles are established, the jet in run t50fm8 should have experienced fiercer expansion than the jet in run t50fm8bp3, so that it is wider at the innermost region, as shown in the upper right panel of Fig. 7. This also causes the thermal and kinetic energy density in the jet in run t50fm8 to be less than that in run t50fm8bp3, as shown in the right panel of Fig. 12. The rapid expansion process might be due to the more prominent hoop effect of the stronger toroidal magnetic field in run t50fm8, as it will make the ejecta interact strongly with the dense gas in front of the jet at the center of the halos. As shown in the lower right panel of Fig. 6, the power of the forward energy outflow in run t50fm8 is actually more dominant than that in run t50fm8bp3 at the innermost region. Then, when the channel is open, the jet in run t50fm8 becomes slimmer than the jet in run t50fm8bp3.
@@
[…]
@@
4.3 Extended phenomena and limitations
The large-scale magnetic field within black hole accretion flows is pivotal for jet launching and collimation (Cao 2011; Chen & Zhang 2021; Li & Cao 2022). Our simulations underscore the criticality of the toroidal magnetic field at kpc-scale in initiating and confining giant radio jets/lobes, as discussed in Sect. 3.2. Additionally, the wake flows of AGN bubbles are essential for elucidating metal-rich outflows and cold filaments in galaxy clusters (Duan & Guo 2018, 2024). This study reveals that jetted bubbles can amplify the weak magnetic fields in the centers of clusters and ultimately result in stronger magnetic fields in their wake flows as shown in Fig. 4 and 9, which may contribute to the seeding of magnetic fields in the circumgalactic and intergalactic media at later stages.
The morphology of GRSs is diverse in observations (Dabhade et al. 2020b), and thus a detailed comparison with our simulations is beyond the scope of this work. Neither do we apply intra-cluster weather effects in our simulations, which explains the bent radio lobes typically seen in clusters (Gan et al. 2017). We observe that many giant radio sources extend along the jet axis (Dabhade et al. 2020b), a feature readily replicated in our simulations, particularly with strong toroidal magnetic field injections (Figs.s 4 and 9). Another notable property is the significant decrease in radio flux at the edges in our simulations (middle panel, Fig. 4), consistent with observations of some large radio lobes (Wu et al. 2020). Furthermore, the gas environment can influence GRS morphology. For instance, in gas halos with lower baryon fractions, jetted lobes may expand more readily, adopting fatter shapes. We aim to delve into the effects of different gas environments in future studies. The velocity will also influence the morphology of radio sources. If higher velocities (for example, fully relativistic) of jets are considered, the density of the jets will be much lower for the same power or kinetic energy flux. This will render the jets less penetrating and produce flatter lobes, as demonstrated in English et al. (2016). On the other hand, full three-dimensional magnetohydrodynamic simulations of jets may encounter kink instability (Mignone et al. 2010) and three-dimensional turbulence, both of which will pose difficulties for the growth of radio sources.
@@ 5 Conclusions
(ii) Behavior of magnetic field. When the injection energy is predominantly toroidal magnetic field compared to thermal or kinetic energy components, the jets tend to produce larger radio sources. Conversely, the application of a stronger poloidal magnetic field during jet injection can effectively suppress the growth of radio lobes.
@@
(iv) Energy transport processes. Higher power jets generally produce jetted bubbles with higher internal pressure, which makes them wider and less penetrating through the centers of clusters, ultimately producing smaller radio sources. At the same time, high-power jets expel energy more rapidly at the centers of galaxy clusters, leaving less energy within the bubbles, thereby impeding the growth of the radio sources. On the contrary, lower power jets can deposit energy more effectively in the outer regions than higher power jets. On the other hand, we also find that energy is mainly transported forward in the inner region and transversely in the outer region for almost all magnetically dominant jets.
@@ Acknowledgements.
This work was supported by the High Performance Computing Center of Henan Normal University. R.Z. was supported by the China Postdoctoral Science Foundation (No. 2023M731014). J.L. was supported by the NSFC (12303020), the Yunnan Fundamental Research Projects (NO.202401CF070169), and the Xingdian Talent Support Plan - Youth Project.
@@ Appendix B Injection of magnetic energy
Using Eqs. 23 - 27, we can analytically solve for $dB_{0}^{2}/dt$, followed by $dB_{\rm inj,r}^{2}/dt$, $dB_{\rm inj,z}^{2}/dt$, and $dB_{\rm inj,\phi}^{2}/dt$. Finally, the magnetic field is updated via the equation:

### v3

@@ 2.3 Jet injection
where $t_{s}$ represents the typical sound crossing time scale in the ambient gas, which is calculated within a characteristic feedback radius $R_{\text{fb}}$. $R_{\text{fb}}$ is the radius within which the initial thermal energy of the ambient gas is equal to the jet outburst energy (Duan & Guo 2020). As $R_{\text{fb}}$  is determined by the initial gas distribution and jet injection energy, both of which are fixed in our simulations, it remains the same as 31 kpc for all our runs. In the context of this work, we have $t_{s}\approx 50\text{ Myr}$ and $P_{\text{fb}}\approx 1.3\times 10^{44}\text{ erg s}^{-1}$. The physical meaning of $P_{\text{fb}}$ lies in the fact that if the jet power, $P_{\text{jet}}$, significantly exceeds $P_{\text{fb}}$, an appreciable fraction of the power deposited by the jet will be dissipated in shocks. Furthermore, taking into account the Eddington luminosity of the black hole $L_{\text{Edd}}\approx 6.71\times 10^{46}\text{ erg s}^{-1}$, the ratio of jet power to Eddington luminosity is $\lambda_{\text{jet}}=P_{\text{jet}}/L_{\text{Edd}}$. For the case where $P_{\text{jet}}=P_{\text{fb}}$, this ratio yields $\lambda_{\text{jet}}\approx 2\times 10^{-3}$.
@@
Figure 2: Scales of the radio sources, $d_{25}$ at the radio power $10^{25}\rm{WHz^{-1}}$ (left) and the corresponding times (right), varying with the jet powers, $P_{\rm jet}$, in simulations with diverse jet parameters, as shown in Table 1. The dots in the right panel represent the time $t_{25}$ when the radio power of the source declines to $10^{25}\rm{WHz^{-1}}$. The horizontal lines in the right panel mark the time $t_{\rm jet}$ when the corresponding jet stops being powered. The triangles denote $t_{\rm jet}+t_{\rm syn}(t_{\rm jet})$ (upward) and $t_{25}+t_{\rm syn}(t_{\rm 25})$ (downward), where $t_{\rm syn}(t)$ is the typical synchrotron cooling time at 144(1+z)MHz for the volume averaged magnetic field strength corresponding to the time $t$, calculated using Eq. 42. The jet powers are normalized by the typical feedback power (bottom) and Eddington luminosity (top), as discussed in Sect. 2.3. The stars are color-coded to represent distinct fractions of energy components and the magnetic structure parameter $\alpha_{p}$, as indicated in the plot.
@@ 3 Results
It is also interesting to see whether the ”lower power-larger bubble” effect exists when considering the same radio power for different sources. In the left panel of Fig. 2, we present the scales of the radio sources varying with jet power, specifically at the instant when the sources have declined to radio power (luminosity) of $10^{25}\rm{WHz^{-1}}$ at 144 (1+z) $\rm{MHz}$ for redshift z=0.3. The methodology for calculating radio emission is detailed in Appendix C. Notably, we can see a trend that the lower power jets result in larger radio sources. The corresponding time $t_{25}$ for the radio power of $10^{25}\rm{WHz^{-1}}$ in different runs are shown in the right panel of Fig. 2 (dots), and we can see they all occur after the jets are switched off ($t_{25}>t_{\rm jet}$ ). We also calculate the typical synchrotron cooling times for the toroidal magnetic field dominant runs at times $t=t_{\rm jet}$ and $t=t_{25}$. We can observe that, if the electrons cannot be adequately accelerated after the jets are switched off, they will quickly cool down, particularly for the high-power jets (upward triangles in the right panel of Fig. 2). This means that, the $t_{25}$ for the high-power jets ($t_{\rm jet}<200$ Myr or $P_{\rm jet}/P_{\rm fb}>0.25$) might be overestimated and when considering synchrotron cooling, the high-power jets will actually result in even smaller radio sources for the same radio power.
@@ 3.3 Radio power of the simulated sources
On the other hand, although jets with significantly lower power can generate much larger radio sources, they may be too faint to observe. As illustrated in Fig. 3, the evolution of the lowest-power jet (run t1200fm8, represented by the black solid line) barely reaches the lower boundary of the observational region. This result implies that the jet powers of these radio sources lie approximately within the range of $10^{-4}-10^{-3}L_{\rm Edd}$ (the corresponding ratios of the jet powers to the Eddington luminosity are shown in Fig. 2). Furthermore, the jet active time of run t1200fm8 is approximately the typical synchrotron cooling time (Eq. 42) of the electrons in the radio lobes, thus the outer part of the lobes might be radio dark if new acceleration of electrons is not adequate. Considering the very low density and pressure of the extreme low-power jets, the mixing between ejecta and surrounding medium might even cause them fail to form low density bubbles or observable radio lobes.
@@ 4.1 Energy transport processes
We examine energy transport processes in different runs before discussing in detail the production of large radio sources. We now focus on two sets of runs: the runs with the same energy fractions ($f_{\rm m}=0.8$, $f_{\rm th}=0.05$, $\alpha_{\rm p}=0.3$) but different powers and the runs with the same power ($t_{\rm jet}=50\ \rm Myr$) but different energy fractions or magnetic field structure. The upper panels in Fig. 5 show the evolution of the distance that the jetted bubble traveled (half of the dimension of radio sources) in these two sets of runs. The evolution of the total energy residing in jetted bubbles is also shown in the lower panels of 5. As an approximation, the bubble region is identified by a density lower than $5\times 10^{-27}\rm{g\ cm^{-3}}$ and a toroidal magnetic field stronger than $10^{-9}\rm{G}$. The dotted vertical lines show when the same total energy is injected in different runs. We can see the scales of the jetted bubbles generally shows positive correlation with the total energy left in the bubbles particularly for the jets with the same energy makeup but different active times (powers). Notably, the high-power jet (run t5fm8) expels energy more rapidly upon jet cessation, leaving less energy within the bubble, which impedes the growth of the radio sources. Actually the most high-power jets can transmit energy outward through more potent shocks compared to low-power jets, which rely on less potent processes such as sound waves and slow expansions (Tang & Churazov 2017; Bambic & Reynolds 2019; Duan & Guo 2020).
While neglecting the mixing processes, the energy outflow rate from the bubble should be equivalent to the power associated with  the work done by the bubble on the surrounding medium via thermal pressure and Maxwell stress (magnetic pressure plus magnetic tension), as shown in Eq. 45. This power can be decomposed into three components: thermal pressure power $P_{out,th}$, magnetic pressure power $P_{out,m1}$, and magnetic tension power $P_{out,m2}$, which can be calculated as follows.
[…]
The evolution of these powers varying with the traveling distance of lobes is shown in Fig. 6. In all cases and in most regions, we can see that the energy transport processes are dominated by either thermal pressure or magnetic pressure. The thermal pressure power and magnetic pressure power in the higher power jet runs are both higher than those in lower power jet runs at the inner regions of clusters, as shown in the left panel. This means that higher power jets experience greater adiabatic losses in the inner regions of clusters, resulting in less energy residing in the bubbles, as discussed in the last paragraph. This should be related to the ”lower power-larger bubble” effect. On the other hand, we can see that lower power jets can transport more energy outside in the outer region of clusters, as indicated by the left panel of Fig. 6. Different energy makeups of jets do not result in as prominent differences as different powers, as shown in the right panel of Fig. 6.
@@ 4.2 Production of large radio sources
To gain some insight into the production of large radio sources, we examine the evolution of the widths of the jetted bubbles and show the results in the upper panels of Fig. 7. We can see that, within the central region of the halos (for example, r¡100kpc), the widths of the bubbles generally exhibit an anti-correlation with the ultimate extent of the simulated radio sources (see Figs. 1 and 5). Furthermore, we check the volume-averaged internal pressure (magnetic pressure plus thermal pressure) within the bubbles and present the results in the lower panels of Fig. 7. We can see that internal pressures of bubbles show a positive correlation with the widths of bubbles in the innermost region of the halos. The dotted vertical lines show where the correlations have just begun to develop. This phenomenon is due to the fact that higher internal pressures within bubbles will cause more prominent transverse expansion, resulting in larger cross sections. The larger cross sections will make the jets much less penetrating and easier to be decelerated by the dense gas in the center of the halos. In the upper left panel of Fig. 7, we can observe that higher power jets generally produce wider bubbles at the center of the halos. Subsequently, there will be stronger interactions between the higher power jets and the surrounding gas, causing them to transfer more energy than lower power jets in the center of the halos, as discussed above.
We then check the different results in run t50fm8 and t50fm8bp3 which are injected with the same energy fractions and power but have different magnetic structures. As the emerging jet is far from equilibrium with the surrounding gas, it will undergo rapid changes in its properties. Before the identified low-density bubbles are established, the jet in run t50fm8 should have experienced fiercer expansion than the jet in run t50fm8bp3, so that it is wider at the innermost region, as shown in the upper right panel of Fig. 7. This also causes the thermal and kinetic energy density in the jet in run t50fm8 to be less than that in run t50fm8bp3, as shown in the right panel of Fig. 9. The rapid expansion process might be due to the more prominent hoop effect of the stronger toroidal magnetic field in run t50fm8, as it will make the ejecta interact strongly with the dense gas in front of the jet at the center of the halos. Then, when the channel is open, the jet in run t50fm8 becomes slimmer than the jet in run t50fm8bp3.
@@
Figure 6: Thermal pressure power (solid lines), magnetic pressure power (dashed lines) and magnetic tension power (dotted lines), calculated using Eqs. 10-12, varying with the traveling distance of lobes. Different colors represent the two sets of runs discussed in Sect. 4.1.
4.3 Extended phenomena and limitations
The large-scale magnetic field within black hole accretion flows is pivotal for jet launching and collimation (Cao 2011; Chen & Zhang 2021; Li & Cao 2022). Our simulations underscore the criticality of the toroidal magnetic field at kpc-scale in initiating and confining giant radio jets/lobes, as discussed in Sect. 3.2. Additionally, the wake flows of AGN bubbles are essential for elucidating metal-rich outflows and cold filaments in galaxy clusters (Duan & Guo 2018, 2024). This study reveals that jetted bubbles can amplify the weak magnetic fields in the centers of clusters and ultimately result in stronger magnetic fields in their wake flows as shown in Fig. 4 and 10, which may contribute to the seeding of magnetic fields in the circumgalactic and intergalactic media at later stages.
The morphology of GRSs is diverse in observations (Dabhade et al. 2020b), and thus a detailed comparison with our simulations is beyond the scope of this work. Neither do we apply intra-cluster weather effects in our simulations, which explains the bent radio lobes typically seen in clusters (Gan et al. 2017). We observe that many giant radio sources extend along the jet axis (Dabhade et al. 2020b), a feature readily replicated in our simulations, particularly with strong toroidal magnetic field injections (Figs.s 4 and 10). Another notable property is the significant decrease in radio flux at the edges in our simulations (middle panel, Fig. 4), consistent with observations of some large radio lobes (Wu et al. 2020). Furthermore, the gas environment can influence GRS morphology. For instance, in gas halos with lower baryon fractions, jetted lobes may expand more readily, adopting fatter shapes. We aim to delve into the effects of different gas environments in future studies. The velocity will also influence the morphology of radio sources. If higher velocities (for example, fully relativistic) of jets are considered, the density of the jets will be much lower for the same power or kinetic energy flux. This will render the jets less penetrating and produce flatter lobes, as demonstrated in English et al. (2016). On the other hand, full three-dimensional magnetohydrodynamic simulations of jets may encounter kink instability (Mignone et al. 2010) and three-dimensional turbulence, both of which will pose difficulties for the growth of radio sources. The turbulence triggered by Kelvin-Helmholtz and Rayleigh-Taylor instabilities in two dimensions cannot experience the energy cascade to small scales as it does in three dimensions, which will result in different entrainment properties and morphologies of the jetted lobes (Massaglia et al. 2016).
@@ 5 Conclusions
(ii) Behavior of magnetic field. When the energy is injected predominantly in the form of toroidal magnetic field, the jets tend to produce larger radio sources. If instead the poloidal field injected is stronger, the growth of the radio lobes is suppressed.
@@
(iv) Energy transport processes. Higher power jets generally produce jetted bubbles with higher internal pressure, which makes them wider and less penetrating through the centers of clusters, ultimately producing smaller radio sources. At the same time, higher power jets expel energy more rapidly at the centers of galaxy clusters through adiabatic losses, resulting from the work done on the surrounding medium by thermal pressure and Maxwell stress. This process leaves less energy within the bubbles produced by higher power jets, thereby impeding the growth of the radio sources.
@@ Acknowledgements.
We thank the anonymous referee for the constructive reports, particularly for the suggestions regarding the addition of discussions. This work was supported by the High Performance Computing Center of Henan Normal University. R.Z. was supported by the China Postdoctoral Science Foundation (No. 2023M731014). J.L. was supported by the NSFC (12303020), the Yunnan Fundamental Research Projects (NO.202401CF070169), and the Xingdian Talent Support Plan - Youth Project.
@@ Appendix B Injection of magnetic energy
Using Eqs. 26 - 30, we can analytically solve for $dB_{0}^{2}/dt$, followed by $dB_{\rm inj,r}^{2}/dt$, $dB_{\rm inj,z}^{2}/dt$, and $dB_{\rm inj,\phi}^{2}/dt$. Finally, the magnetic field is updated via the equation: